OpenStudy (anonymous):

What is the value of the expression? 80 + 4^3 – 6 · 5 A. 62 B. 114 C. 430 D. 690

OpenStudy (anonymous):

43 is 4^3

OpenStudy (anonymous):

the power is 3 over 4

OpenStudy (anonymous):

Easy way to determine order of operations; PFMDAS= Please Forgive Dear My Aunt Sally= Perentheses, Fractions, Division, Multiplication, Addition, Subtraction. 80+64= 144 144-30= 114 The answer is 114.

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

@mango45

OpenStudy (anonymous):

with

OpenStudy (anonymous):

@thecacaman822

OpenStudy (anonymous):

wat

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

@harz360

OpenStudy (anonymous):

@love14

OpenStudy (anonymous):

i need help

OpenStudy (anonymous):

B. 114

OpenStudy (anonymous):

thank you love14

OpenStudy (anonymous):

80+ 4 ^3 =144 - 6 *5 =11

OpenStudy (anonymous):

114***

OpenStudy (anonymous):

80 + 4^3 – 6 · 5 80 + 64 - 6 · 5 80 + 64 - 30 144 - 30 = 114 PEMDAS P- Parentheses E- Exponent M- Multiply D- Divide A- Add S- Subtract The answer is B.

OpenStudy (whpalmer4):

\[80+4^3-6*5\]no parentheses, so do the exponent: \[80+64-6*5\]next do multiplication \[80+64-30\]finally do addition and subtraction \[144-30 = 114\]